Ladder AB is leaning against building AC. The foot of the leader is a distance CB from the building. AC = 25ft. And CB= 12ft
A) find the length of the ladder to the nearest foot. Explain how you arrived at your answer. Please help me algebra homework due tomorrow

Answer by checkley79(3341)

12^2+25^2=C^2
144+625=C^2
769=C^2
C=SQRT769
C=28 FT. IS THE LENGTH OF THE LADDER.